A student group is producing activity calendars for sale at the beginning of the academic year and must decide how many to have printed. The calendars cost $2 each to print and sell for $5. Those not sold have no salvage value. Demand is normally distributed with a mean of 3000 and standard deviation of 100.

Cs = Sales price/unit – Cost/unit  Co = Cost/unit – Salvage value  Service Level =
Q = µ + Z σ

  1. What is the optimal service level?
  2. How many calendars should the organization produce?
Answer #1

COST PRICE = 2
SALES PRICE = 5
SALVAGE VALUE = 0

Cost of shortage(Cs) = sales price - cost price = 5 - 2 = 3
Cost of overage(Co) = Cost price - salvage value = 2 - 0 = 2

Service level = Cs / (Cs + Co) = 3 / (3 + 2) = 0.6
The Z value that corresponds to a service level of 0.6 is = 0.25

Mean demand = 3000
Standard deviation = 100

Optimal stocking level = Mean demand + (Z * Standard deviation)

= 3000 + (0.25 * 100) = 3025

A Thumbs Up! would be really helpful for me. If you have any questions, please leave a comment and I will get back to you as soon as possible.